McIver's Meals, Inc. currently pays a $1.00 annual dividend. Investors believe that dividends will grow at 15% next year, 10% annually for the two years after that, and 5% annually thereafter. Assume the required return is 10%. What is the current market price of the stock?
Answer #1

D1=(1*1.15)=1.15

D2=(1.15*1.1)=1.265

D3=(1.265*1.1)=1.3915

Value after year 3=(D3*Growth Rate)/(Required rate-Growth Rate)

=(1.3915*1.05)/(0.1-0.05)

=$29.2215

Hence current price=Future dividends and value*Present value of discounting factor(rate%,time period)

=1.15/1.1+1.265/1.1^2+1.3915/1.1^3+29.2215/1.1^3

=$25.09(Approx).